2. A 60-watt light bulb has a voltage of 120 volts applied across it and a current of 0.5 A flows through the bulb. What is the resistance of the light bulb?




3. A resistance of 60 ohms has 0.4 A of current flowing when it is connected across a battery. What is the voltage of the battery?




4. The resistance of a small electric motor is 24 ohms. If the motor operates properly on a voltage of 6 volts, what current does the motor require?



1
Expert's answer
2021-04-06T18:24:06-0400

2)


"R=\frac{V}{I}=\frac{120}{0.5}=240\Omega"

3)


"V=IR=(0.4)(60)=24\ V"

4)


"I=\frac{V}{R}=\frac{6}{24}=0.25\ A"